# ORB-SfMLearner: ORB-Guided Self-supervised Visual Odometry with Selective Online Adaptation

### v1

@@ Abstract
Deep visual odometry, despite extensive research, still faces limitations in accuracy and generalizability that prevent its broader application. To address these challenges, we propose an Oriented FAST and Rotated BRIEF (ORB)-guided visual odometry with selective online adaptation named ORB-SfMLearner. We present a novel use of ORB features for learning-based ego-motion estimation, leading to more robust and accurate results. We also introduce the cross-attention mechanism to enhance the explainability of PoseNet and have revealed that driving direction of the vehicle can be explained through attention weights, marking a novel exploration in this area. To improve generalizability, our selective online adaptation allows the network to rapidly and selectively adjust to the optimal parameters across different domains. Experimental results on KITTI and vKITTI datasets show that our method outperforms previous state-of-the-art deep visual odometry methods in terms of ego-motion accuracy and generalizability.
@@ I Introduction
Estimating camera ego-motion from monocular videos is essential for various computer vision and robotics tasks. Especially recently, concentrated research [1, 2] on 3D representation has led to higher demands for camera pose estimation. Traditional methods [3, 4, 5, 6] find matches across frames and restore camera transformations with epipolar geometry, while learning-based self-supervised methods [7, 8, 9] usually infer depth and ego-motion simultaneously and then establish a self-supervised constraint with the photometric reconstruction error. Learning-based methods [10, 11] have been widely studied in recent years due to its fast inference and ability to learn high-level features from data.
@@
To address these problems, several works have been proposed. Training on larger datasets may help mitigate effect of domain gaps and achieve better accuracy. Wang et al. [12] use synthetic data to attempt large-scale VO training. However, due to the complexity of real-world environments, it’s challenging to gather a sufficient amount of data, which also requires a significantly longer training time. The self-supervised nature of current learning-based methods provides another solution. Some recent works [13, 14] focus on online fine-tuning of pre-trained VO models during test time. This learning while testing approach proves to be very effective. However, in scenarios where directly training on the entire test set fails to yield satisfactory results, online fine-tuning hardly works well. Therefore, more robust training strategies are still needed to allow the model to achieve better performance independently.
@@
  • To enhance interpretability of the networks’ learning process, we intuitively explored the impact of ORB features. As one of the earliest works to explore the interpretability within the PoseNet, we aim to provide insights for related research.
@@ III-E Selective Online Adaptation
After pre-training the PoseNet and DepthNet on the source domain, we continue to adapt the model during inference with our proposed selective online adaptation (SOA), as illustrated in Algorithm 1. We first prepare the incoming frames with the proposed ORB-Augmentation. Then we conduct inference with current PoseNet and DepthNet and use the estimated relative pose $P$ and depth $D_{i},D_{i+1}$ to perform differentiable warping (Section III-B). Here, the loss from this self-supervised process is used to update the model parameters. However, we don’t directly iterate to optimize and output the results. Instead, we infer with the updated model to compute the loss again, and only if the loss decreases do we select the updated parameters. We call this strategy selective adaptation. For simplicity, Algorithm 1 only shows the case where each input snippet consists of 2 frames. However, we can also input more frames and perform differentiable warping on each pair of frames, updating the model online using the average loss.

### v2

@@ Abstract
Deep visual odometry, despite extensive research, still faces limitations in accuracy and generalizability that prevent its broader application. To address these challenges, we propose an Oriented FAST and Rotated BRIEF (ORB)-guided visual odometry with selective online adaptation named ORB-SfMLearner. We present a novel use of ORB features for learning-based ego-motion estimation, leading to more robust and accurate results. We also introduce the cross-attention mechanism to enhance the explainability of PoseNet and have revealed that driving direction of the vehicle can be explained through the attention weights. To improve generalizability, our selective online adaptation allows the network to rapidly and selectively adjust to the optimal parameters across different domains. Experimental results on KITTI and vKITTI datasets show that our method outperforms previous state-of-the-art deep visual odometry methods in terms of ego-motion accuracy and generalizability.
@@ I Introduction
Estimating camera ego-motion from monocular videos is essential for various computer vision and robotics tasks. Especially recently, concentrated research [1, 2] on 3D representation has led to higher demands for camera pose estimation. Traditional methods [3, 4, 5, 6] find matches across frames and restore camera transformations with epipolar geometry, while learning-based self-supervised methods [7, 8, 9] usually infer depth and ego-motion simultaneously and then establish a self-supervised constraint with the photometric reconstruction error. Learning-based methods have been widely studied in recent years due to its fast inference and ability to learn high-level features from data [10, 11].
@@
To address these problems, several works have been proposed. Training on larger datasets may help mitigate effect of domain gaps and achieve better accuracy. Wang et al. [12] use synthetic data to attempt large-scale VO training. However, due to the complexity of real-world environments, it’s challenging to gather a sufficient amount of data, which also requires a significantly longer training time. The self-supervised nature of current learning-based methods provides another solution. Some recent works [13, 14] focus on online fine-tuning of pre-trained VO models during test time. This learning while testing approach proves to be very effective. However, in scenarios where directly training on the entire test set fails to yield satisfactory results, online fine-tuning hardly works well. Therefore, more robust training strategies are still needed for the model to achieve better performance independently.
@@
  • To enhance interpretability of the networks’ learning process, we intuitively explored the impact of ORB features. As one of the earliest works to explore the interpretability of ego-motion estimation, we aim to provide insights for related research.
@@ III-E Selective Online Adaptation
After pre-training the PoseNet and DepthNet on the source domain, we continue to adapt the model during inference with our proposed selective online adaptation (SOA), as illustrated in Algorithm 1. We first prepare the incoming frames with the proposed ORB-Augmentation. Then we conduct inference with current PoseNet and DepthNet and use the estimated relative pose $P$ and depth $D_{i},D_{i+1}$ to perform differentiable warping (Section III-B). Here, the loss from this self-supervised process is used to update the model parameters. However, we don’t directly iterate to optimize and output the results. Instead, we infer with the updated model to compute the loss again, and only if the loss decreases do we select the updated parameters. We call this strategy selective adaptation. For simplicity, Algorithm 1 only shows the case where each input snippet consists of 2 frames. We can also input more frames and perform differentiable warping on each pair of frames, updating the model online using the average loss.